13. The data table below shows the number of computers sold at the Bronco Store in a week, based on online advertising. (30 points possible).

Online Ads               Computers Sold

2                               25       

1                               10

4                               30

1                               10

Sx = 10

2                               25

Sy = 100

a. Calculate the means for both Ads and Computers Sold.

              

b. Calculate the slope for the estimated regression equation.

The (x-x)(y-y) = 40, just need to calculate denominator.
Answer #1

a)

here,

means for online ads x̅ = Σx / n= 10/5 = 2 ,

mean for somputers sold ȳ = Σy/n = 100/5 = 20

b)

x y (x-x̅)²
2 25 0.00
1 10 1.00
4 30 4.00
1 10 1.00
2 25 0.00

so, denominator = Σ(x-x̅)² =    6.0000

Slope = (x-x)(y-y)/Σ(x-x̅)² = 40/6 = 6.6667